OpenStudy (anonymous):

Construct ecamples. Let f and g be functions that are neither even not odd. a. create an example where f+g is even b. create an example where f+g is odd c. create an example where f times is even d. create an example where f times is odd

OpenStudy (michele_laino):

question a) I consider this pair of functions: \[\begin{gathered} f\left( x \right) = {x^3} - 1 \hfill \\ g\left( x \right) = - {x^3} + {x^2} + 1 \hfill \\ \end{gathered} \]